OpenStudy (anonymous):

if a plane can travel 450 miles per hour with the wind and 410 miles per hour against the wind, find the speed of the plane without a wind and speed of the wind? ***i know the speed is 230 but i cant get the wind speed HELP****

OpenStudy (anonymous):

When you are flying with the wind, it adds to your speed, and when you are flying against it, it takes away from it. Let's call the Speed of the Plane 'x' and the Speed of the Wind 'y'. Therefore we know that x + y = 450 and x - y = 410

OpenStudy (anonymous):

You then get either x or y by itself, plug in and simply to find the answer. For example: x - y = 410 x = y + 410 and then we plug that into x + y = 450 to get x + y = 450 (y + 410) + y = 450 2y + 410 = 450 2y = 40 y = 20

OpenStudy (anonymous):

Remembering that we found x to be equal to (y + 410) can now use the value of y to find the value of x. x = y + 410 y = 20 x = 20 + 410 x = 430 So x is 430 and y is 20, therefore the Speed of the Plane is 430 and the Speed of the Wind is 20

OpenStudy (anonymous):

The average speed, the speed without the wind is the sum of (450+410)/2 or 430mph. The wind speed is 450-430 or 20mph.